OpenStudy (anonymous):

write the repeating decimal as a fraction of 0.7 and 0.18

OpenStudy (anonymous):

what r the steps to getting the answer?

Parth (parthkohli):

Use long division to divide 70 by 18.

OpenStudy (anonymous):

btw they r both seperate answers so i have to make the repeating decimal as fraction of of 0.7 first

Parth (parthkohli):

...what?

OpenStudy (anonymous):

ok so the question is: write the rpeating decimal as a fraction: 0.7

OpenStudy (anonymous):

and also of 0.18

OpenStudy (anonymous):

so i have to write the 0.7 as a fraction first

Parth (parthkohli):

0.7 = 7/10 and 0.18 = 18/100. That one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank u :)